OpenStudy (anonymous):

factor 12x^2 - 72x + 77 = 0

OpenStudy (anonymous):

use quadratics

OpenStudy (anonymous):

what would the answer be though?

OpenStudy (anonymous):

x=4.6072751268321595 x=1.3927248731678408

OpenStudy (anonymous):

thank you. Are you good with calc because I have one other calc question

OpenStudy (anonymous):

lol idk i can try

OpenStudy (anonymous):

x=16(18+93−−√),16(18−93−−√) x≈4.607275126832,1.392724873168 Yep he's right

OpenStudy (anonymous):

Assume that the demand for tuna in a small coastal town is given by p = 600,000/q^1.5 where q is the number of pounds of tuna that can be sold in a month at p dollars per pound. Assume that the town's fishery wishes to sell at least 5,000 pounds of tuna per month. (a) How much should the town's fishery charge for tuna in order to maximize monthly revenue? HINT [See Example 3, and don't neglect endpoints.] (Round your answer to the nearest cent.) p = $ per lb (b) How much tuna will it sell per month at that price? q = lb (c) What will be its resulting revenue? (Round your answer to the nearest dollar.) $ per month